a rectangle has a perimeter of 1000 feet and the length is 168 feet more than the width. what is the width and the length

Answer by Cromlix(4381)

Width = x
Length = x + 168
Perimeter = 2*Width + 2*Length
1000 = 2*x + 2*(x + 168)
1000 = 2x + 2x + 336
4x = 1000 - 336
4x = 664
x = 166
Width = 166 feet
Length = 334 feet.
Hope this helps.
:-)